OpenStudy (anonymous):

What is the slope of the line 5x - 2y = -10? Please help. :O

OpenStudy (e.cociuba):

y=5/2x + 5

OpenStudy (anonymous):

My options for answers are: -5/2 5/2 2/5 -2/5 If you could show me how you got it, that'd be great.

OpenStudy (e.cociuba):

5/2

OpenStudy (anonymous):

the slope is 5/2. What e.cocuiba did was this: 5x-2y = -10. Subtract 5x from both sides giving -2y = -10 - 5x. Divide both sides by -2 giving you your final equation: y=5+5/2x

OpenStudy (anonymous):

Thanks so much :)

OpenStudy (e.cociuba):

np:) Thanks @brandi2013 for explaining